OpenStudy (anonymous):

Alright. And so the problem is a little long, it's an exponential problem. I usually get these but this one is different. The world population in 2005 was 6.2 billion and growing exponentially at a rate of 1.14% per year. the function P(t)=6.2(10^0.005t) provides a good model for the population growth pattern. In this model what does P(t) represent and what does t represent Also says Explain how you can be sure that this function givesthe correct popuation in 2005

OpenStudy (wolf1728):

If the population increases by 1.14% per year shouldn't the formula be P(t)=6.2*(10^0.0114t) and NOT P(t)=6.2(10^0.005t) P(t) represents the world's population (in billions) after 't' years. It gives the correct population for 2005 because P(t)=6.2*(10^0.0114t) t =0 and so P(t) = 6.2 * 10^0 = 6,2 and that is the population stated already.

OpenStudy (anonymous):

Thanks this really helped. Yea I think it should be .0114 it's probably just a typo since that's what my paper says.

OpenStudy (wolf1728):

Population after 1 year P(t)=6.2*(10^0.0114*1) P(t)=6.2*(10^0.0114*1) P(t)=6.2*10^.0114 P(t)=6.2*10^1.0265970218 P(t)=6.508 billion

OpenStudy (wolf1728):

okay glad to help out

OpenStudy (anonymous):

This was seriously really helpful! Thanks A LOT!